# Efficient Best-of-Both-Worlds Algorithms for Contextual Combinatorial Semi-Bandits

### v1

@@ Abstract
We introduce the first best-of-both-worlds algorithm for contextual combinatorial semi-bandits that simultaneously guarantees $\widetilde{\mathcal{O}}(\sqrt{T})$ regret in the adversarial regime and $\widetilde{\mathcal{O}}(\ln T)$ regret in the corrupted stochastic regime. Our approach builds on the Follow-the-Regularized-Leader (FTRL) framework equipped with a Shannon entropy regularizer, yielding a flexible method that admits efficient implementations. Beyond regret bounds, we tackle the practical bottleneck in FTRL (or, equivalently, Online Stochastic Mirror Descent) arising from the high-dimensional projection step encountered in each round of interaction. By leveraging the Karush-Kuhn-Tucker conditions, we transform the $K$-dimensional convex projection problem into a single-variable root-finding problem, dramatically accelerating each round. Empirical evaluations demonstrate that this combined strategy not only attains the attractive regret bounds of best-of-both-worlds algorithms but also delivers substantial per-round speed-ups, making it well-suited for large-scale, real-time applications.
@@ 1 Introduction
Online decision-making is often modelled via the multi-armed bandit framework: over $T$ rounds, a learner selects an action and incurs a loss, observing only partial feedback. Many real-world tasks—from selecting at most $m$ movies on a streaming platform to curating a list of $m$ products on an e-commerce homepage—require choosing a subset of up to $m$ items from $K\gg m$ base arms in each round. This combinatorial bandit variant admits either semi-bandit feedback (per-arm losses) or full-bandit feedback (aggregate loss). Semi-bandit feedback is realistic—web analytics record click-through outcomes for each displayed item—and statistically advantageous, reducing minimax regret from $\mathcal{O}(\sqrt{KT})$ under full-bandit feedback to $\mathcal{O}(\sqrt{mT})$ (audibert2014regret, ). In each round, the learner faces either stochastic losses (e.g., from a fixed linear model on random contexts) or adversarial losses (modeling malicious perturbations). Contexts consist of user-feature vectors drawn i.i.d. from a distribution, and instantaneous losses are linear in these features. The best-of-both-worlds (BOBW) paradigm unifies strategies across stochastic and adversarial loss regimes, achieving optimal regret guarantees in both (zimmert2019beating, ; tsuchiya2023further, ). However, gaps remain in current BOBW approaches for contextual combinatorial bandits—highlighted by tsuchiya2023further ; kuroki2024best and exacerbated by evolving regulations and emerging technologies—motivating us to revisit the problem.
@@
  • Adversarial and corrupted stochastic regimes. Advertising markets, for instance, evolve in real time as competitors react to each other (balseiro2015repeated, ; jin2018real, ), while recommendation systems must adapt to shifting user preferences (koren2009collaborative, ) and defend against malicious actors (mukherjee2013spotting, ; christakopoulou2019adversarial, ; zhang2020practical, ). A natural remedy is a best-of-both-worlds strategy—provided it is computationally efficient. However, recent privacy regulations (e.g., the EU’s GDPR (GDPR2016, )) restrict context collection (for example, via third-party cookies), so the assumption of perfect knowledge of the context distribution no longer holds. This gap motivates BOBW algorithms that explicitly handle corrupted stochastic contexts—treating unexpected losses as corruptions—to secure optimal regret guarantees, which existing literature has yet to address.
  • Subroutine computational efficiency for combinatorial bandits. Advances in large language models (LLMs) allow platforms to generate vast pools of $K$ candidate ads at negligible cost, which can then be deployed for personalized advertising (meguellati2024good, ). Yet most existing algorithms require solving a $K$-dimensional convex program in each round, making them increasingly expensive as $K$ grows. Moreover, on-the-fly customization of ad attributes (e.g., color, layout) via generative models further increases the runtime of these subroutines. Consequently, accelerating the per-round projection step is essential for practical, large-scale deployment.
Contributions. We design an algorithm that provably addresses the stated challenges and present the following two key contributions:
  • Best-of-Both-Worlds for Contextual Combinatorial Semi-Bandits. We achieve both $\widetilde{\mathcal{O}}(\sqrt{T})$ regret in the adversarial regime and $\widetilde{\mathcal{O}}(\ln T)$ regret in the corrupted stochastic regime. Our method instantiates Follow-the-Regularized-Leader (FTRL) with a Shannon-entropy regularizer, enabling efficient implementations in many practical settings (see Sec. 2).
  • Accelerated Projection for FTRL/OSMD. We accelerate the projection subroutine in any FTRL scheme with a Legendre regularizer—equivalently, in Online Stochastic Mirror Descent (OSMD) for linear payoffs under the corresponding mirror map (shalev2012online, ). By exploiting the Karush-Kuhn-Tucker (KKT) conditions, we reduce the typical $K$-dimensional convex projection to a one-dimensional root-finding problem (see Sec. 3).
As a result, our algorithm is comparable to the per-iteration complexity of Follow-the-Perturbed-Leader (FTPL) (neu2015first, ; neu2016importance, )—which injects random noise to cumulative losses and selects the action with minimal perturbed total loss—while preserving the tight adversarial and stochastic regret guarantees characteristic of FTRL. Hence, we achieve both statistical and computational efficiency.
@@ 1.1 Related Work
The contextual combinatorial bandit problem was introduced by Qin et al. qin2014contextual , who proposed the C2UCB algorithm for multi-item recommendations. This framework builds on earlier combinatorial semi-bandit models motivated by influence maximization in social networks chen2013combinatorial . Under the i.i.d. reward assumption, linear generalization across arms reduces the regret dependence on the action-set size $K$ from $\sqrt{K}$ to the context dimension $d$. Consequently, when $d\ll K$, the regret scales with $\sqrt{dT}$ instead of $\sqrt{KT}$, offering a substantial statistical advantage. Stochastic variants have explored Thompson sampling wang2018thompson and refined confidence sets takemura2021near ; however, none of these algorithms provide optimal regret guarantees against an adaptive adversary.
(ii) Semi-Bandit Feedback and Efficient Optimization.
@@
Combinatorial bandits were first formalized by cesa2012combinatorial , who generalized Exp3 (auer2002nonstochastic, ) from single arms to binary action vectors. Their algorithm, ComBand, operates under full-bandit feedback—the learner only observes an aggregated loss for the binary action vector they have chosen—and achieves $\widetilde{\mathcal{O}}\bigl{(}\sqrt{mKT\ln(K/m)}\bigr{)}$ regret. Follow-up work connected this update to mirror descent on combinatorial polytopes: ComponentHedge koolen2010hedging , CombEXP combes2015combinatorial , and the OSMD analysis of audibert2014regret . Bubeck and Cesa-Bianchi bubeck2012regret provide a concise taxonomy and prove matching lower bounds, showing that the $\sqrt{mKT}$ rate is information-theoretically optimal whenever semi-bandit feedback is available. To reduce per-round runtime, Neu and Bartók neu2016importance introduced a stochastic mirror-descent update that samples a single action and updates only the observed coordinates, thereby avoiding the full Bregman projection required by standard OSMD or FTRL. The trade-off is an extra logarithmic factor in the regret bound. Variance-reduced estimators and high-probability analyses later refined these guarantees (zimmert2019beating, ), yet computing the exact Bregman projection still requires time linear in the number of arms $K$.
(iii) Adversarial Linear Bandits and Best-of-Both-Worlds.
@@
When losses can adapt to the learner’s past, the i.i.d. assumption no longer holds. For single-arm actions ($m=1$) the classical EXP4 algorithm attains $\widetilde{\mathcal{O}}(\sqrt{T})$ regret, but its running time and memory scale with the number of experts—that is, the total number of deterministic policies mapping contexts to arms—which grows exponentially in the arm size $K$ (auer2002nonstochastic, ). Neu and Olkhovskaya neu2020efficient mitigated this explosion with RealLinExp3, achieving $\widetilde{\mathcal{O}}\sqrt{dKT}$ regret under the strong assumption that the context distribution is known. liu2023bypassing proposed an algorithm that achieves $\mathcal{O}((dm)^{3}\sqrt{T})$ regret bound when applied to combinatorial semi-bandit case with only access to $1$ context sample per time period $t$, though the resulting log-determinant FTRL remains intractable for the combinatorial bandits case because of exponentially many constraints foster2020adapting ; zimmert2022return and the nonlinearity introduced by their lifting covariance technique. The first near-optimal treatment is the Matrix-Geometric-Resampling (MGR) algorithm of Zierahn et al. zierahn2023nonstochastic , which achieves $\widetilde{\mathcal{O}}\bigl{(}\sqrt{mKT\max\{d,m/\lambda_{\text{min}}(\Sigma)\}}\bigr{)}$ regret but fails to provide best-of-both-worlds regret guarantees and relies on a sampling sub-routine which requires $\mathcal{O}(\ln T)$ samples in each round. Similar to the issue of generalizing (liu2023bypassing, ), a straightforward extension of the best-of-both-worlds (BOBW) algorithm from the linear contextual bandit setting, as studied by kuroki2024best , to the contextual combinatorial bandit framework yields an action space whose cardinality grows exponentially with the size of the combinatorial decision variables. This exponential growth renders the resulting optimization problem NP-hard and computationally infeasible for large-scale applications. More broadly, the BOBW question—achieving $\widetilde{\mathcal{O}}(\ln T)$ under stochastic contexts and $\widetilde{\mathcal{O}}(\sqrt{T})$ under adversarial ones without prior knowledge—originated in the $K$-armed bandit setting (bubeck2012best, ; seldin2014one, ) and has since been settled for linear bandits via data-dependent stability (lee2021achieving, ), and a streamlined FTRL scheme (kong2023best, ) and for contextual bandits via Exp4-style expert reductions (pacchiano2022best, ; dann2023blackbox, ), at the cost of exponential policy-enumeration. In the purely combinatorial semi-bandit setting, hybrid-regularizer methods (zimmert2019beating, ; ito2021hybrid, ) achieve the optimal BOBW rates but do not consider contextual information and the stemmed difficulty of inverse-covariance matrix estimation. Consequently, no prior method simultaneously handles unknown covariance, large/infinite policy classes, and combinatorial action structure while maintaining BOBW regret bounds. We close this gap with a Shannon-entropy FTRL algorithm that requires neither policy enumeration nor known $\Sigma$, yet still achieves optimal BOBW rates on arbitrary combinatorial action sets.
@@ 2 Best-of-Both-Worlds Algorithm for Contextual Combinatorial Semi-Bandits
We begin our analysis of contextual combinatorial bandits by describing the interaction protocol that the learner follows and the problem settings that the best-of-both-worlds framework unifies. Given $K$ base arms, $m$ maximum number of base arms allowed to pull per round, an action set $\mathcal{A}=\{A\in\{0,1\}^{K}:\sum_{k=1}^{K}(A)_{k}\leq m\}$, and a context space $\mathcal{X}\subset\mathbb{R}^{d}$, the interaction protocol for the contextual combinatorial bandit problem proceeds as follows.
@@
In each round $t=1,\dots,T$ the interaction protocol proceeds as follows. The environment first chooses the loss coefficients $\theta_{t,1},\ldots,\theta_{t,K}\in\mathbb{R}^{d}$, and draws an i.i.d. context $X_{t}\sim\mathcal{D}$. The learner observes $X_{t}$ and plays the vector $A_{t}\in\mathcal{A}$. The learner then observes the arm-wise losses $\ell_{t}(X_{t},k)=\langle X_{t},\theta_{t,k}\rangle$ for every $k$ such that $(A_{t})_{k}=1$ and suffers the total loss $\sum_{k=1}^{K}\ell_{t}(X_{t},k)(A_{t})_{k}$. Consistent with the prior work on linear contextual bandits such as (kuroki2024best, ), we adopt the following assumption throughout this paper.
@@
Under this assumption, we distinguish between the adversarial and stochastic regimes as follows. In the adversarial regime, the unknown loss coefficients $\{\theta_{t,k}\}_{k=1}^{K}$ may vary arbitrarily with respect to time but is independent of the learner’s action sequence. In the stochastic regime, the loss function takes the form $\ell_{t}(X_{t},k)=\langle X_{t},\theta_{k}\rangle+\varepsilon(X_{t},k)$ for all $k\in[K]$, where $\theta_{k}$ is also unknown but fixed throughout the $T$ interaction periods, and $\varepsilon(X_{t},k)$ is an independent and bounded zero-mean noise. There is an additional intermediate regime that interpolates between the adversarial and stochastic regime, referred to as corrupted stochastic regime (zimmert2021tsallis, ; kuroki2024best, ). In this case, the loss function associated with each base arm $k$ is defined by $\ell_{t}(X_{t},k)=\langle X_{t},\theta_{t,k}\rangle+\varepsilon(X_{t},k)$, where $\varepsilon(X_{t},k)$ is independent and bounded zero-mean noise. On the other hand, the coefficients $\theta_{t,k}$ are such that there exist fixed and unknown vectors $\theta_{1},\ldots,\theta_{K}$ and that satisfy $\sum_{t=1}^{T}\max_{A\in\mathcal{A}}\sum_{k=1}^{K}\|\theta_{t,k}-\theta_{k}\|_{2}(A)_{k}\leq C$ for a fixed corruption level constant $C>0$. Note that $C=0$ corresponds to the stochastic regime and $C=\Theta(T)$ corresponds to the adversarial regime with additional zero-mean noise. For any fixed context vector $x\in\mathcal{X}$, we define $u^{\star}:\mathcal{X}\to\mathcal{A}$ as the optimal context-dependent action map that achieves minimum loss in hindsight, that is,
@@
where the expectation is taken with respect to the action-selection distribution chosen by the learner and the sequence of random contexts and loss coefficients chosen by the environment. As noted in the introduction, it is well-established in the literature that the optimal regret bounds are $\widetilde{\mathcal{O}}(\ln T)$ in the stochastic regime and $\widetilde{\mathcal{O}}(\sqrt{T})$ in the adversarial regime. Algorithms that achieve both rates simultaneously and without prior knowledge of the environment’s nature are referred to as “best-of-both-worlds” (bubeck2012best, ; seldin2014one, ).
@@
4:  Compute $\bar{A}_{t}(X_{t})\in\operatorname*{argmin}_{a\in\operatorname*{conv}(\mathcal{A})}\sum_{s=1}^{t-1}\sum_{k=1}^{K}\langle X_{s},\tilde{\theta}_{s,k}\rangle(\bar{A})_{k}+\psi_{t}(a)$
@@
12:   end for
13:   return $\widehat{\Sigma}_{t,k}^{+}=(I+\sum_{n=1}^{M_{t}}C_{n,k})/2$ for all $k\in[K]$
14:  End Subroutine
15:  Compute $\tilde{\theta}_{t,k}=\widehat{\Sigma}_{t,k}^{+}X_{t}\ell_{t}(X_{t},k)(A_{t})_{k}$ for all $k\in[K]$
16:end for

### v2

@@ Abstract
We introduce the first best-of-both-worlds algorithm for contextual combinatorial bandits that simultaneously guarantees $\widetilde{\mathcal{O}}(\sqrt{T})$ regret in the adversarial regime and $\widetilde{\mathcal{O}}(\ln T)$ regret in the corrupted stochastic regime. Our approach builds on the Follow-the-Regularized-Leader (FTRL) framework equipped with a negative Shannon entropy regularizer, yielding a flexible method that admits efficient implementations. Beyond regret bounds, we tackle the practical bottleneck in FTRL (or, equivalently, Online Stochastic Mirror Descent) arising from the high-dimensional projection step encountered in each round of interaction. By leveraging the Karush-Kuhn-Tucker conditions, we transform the $K$-dimensional convex projection problem into a single-variable root-finding problem, dramatically accelerating each round. Empirical evaluations demonstrate that this combined strategy not only attains the attractive regret bounds of best-of-both-worlds algorithms but also delivers substantial per-round speed-ups, making it well-suited for large-scale, real-time applications.
@@ 1 Introduction
Online decision-making is often modeled via the multi-armed bandit framework: over $T$ rounds, a learner selects an action and incurs a loss, observing only partial feedback. Many real-world tasks—from selecting at most $m$ movies on a streaming platform to curating a list of $m$ products on an e-commerce homepage—require choosing a subset of up to $m$ items from $K\gg m$ base arms in each round. This combinatorial bandit variant admits either semi-bandit feedback (per-arm losses) or full-bandit feedback (aggregate loss). Semi-bandit feedback is realistic—web analytics record click-through outcomes for each displayed item—and statistically advantageous: for $m$-set combinatorial bandits it improves the minimax rate from $\widetilde{\mathcal{O}}(m\sqrt{KT})$ under full-bandit feedback to $\widetilde{\mathcal{O}}(\sqrt{mKT})$ under semi-bandit feedback (Audibert et al., 2014). In each round, the learner faces either stochastic losses (e.g., from a fixed linear model on random contexts) or adversarial losses (modeling malicious perturbations). Contexts consist of user-feature vectors drawn i.i.d. from a distribution, and instantaneous losses are linear in these features. The best-of-both-worlds (BOBW) paradigm unifies strategies across stochastic and adversarial loss regimes, achieving optimal regret guarantees in both (Zimmert et al., 2019; Tsuchiya et al., 2023). However, gaps remain in current BOBW approaches for contextual combinatorial bandits—highlighted by Tsuchiya et al. (2023); Kuroki et al. (2024) and exacerbated by evolving regulations and emerging technologies—motivating us to revisit the problem.
@@
  • Adversarial and Corrupted Stochastic Regimes. Advertising markets, for instance, evolve in real time as competitors react to each other (Balseiro et al., 2015; Jin et al., 2018), while recommendation systems must adapt to shifting user preferences (Koren, 2009) and defend against malicious actors (Mukherjee et al., 2013; Christakopoulou and Banerjee, 2019; Zhang et al., 2020). A natural remedy is a best-of-both-worlds strategy—provided it is computationally efficient. However, recent privacy regulations (e.g., the EU’s GDPR (GDPR, 2016)) restrict context collection (for example, via third-party cookies), so the assumption of perfect knowledge of the context distribution no longer holds. This gap motivates BOBW algorithms that explicitly handle corrupted stochastic contexts—treating unexpected losses as corruptions—to secure optimal regret guarantees, which existing literature has yet to address.
  • Subroutine Computational Efficiency for Combinatorial Bandits. Advances in large language models (LLMs) allow platforms to generate vast pools of $K$ candidate ads at negligible cost, which can then be deployed for personalized advertising (Meguellati et al., 2024). Yet most existing algorithms require solving a $K$-dimensional convex program in each round, making them increasingly expensive as $K$ grows. Moreover, on-the-fly customization of ad attributes (e.g., color, layout) via generative models further increases the runtime of these subroutines. Consequently, accelerating the per-round projection step is essential for practical, large-scale deployment.
Contributions. Our key contributions are summarized as follows.
[…]
  • Accelerated Projection for FTRL/OSMD. For a popular class of combinatorial action sets, namely the $m$-set, we accelerate the projection subroutine in any FTRL scheme with a Legendre regularizer—equivalently, in Online Stochastic Mirror Descent (OSMD) for linear payoffs under the corresponding mirror map (Shalev-Shwartz, 2012). By exploiting the Karush-Kuhn-Tucker (KKT) conditions, we reduce the typical $K$-dimensional convex projection to a one-dimensional root-finding problem (see Sec. 3). Results therein are generalizable to a broader class of combinatorial action sets that admit a separable structure, e.g., separable matroids.
As a result, our algorithm is comparable to the per-iteration complexity of Follow-the-Perturbed-Leader (FTPL) (Neu, 2015; Neu and Bartók, 2016)—which injects random noise to cumulative losses and selects the action with minimal perturbed total loss—while preserving the tight adversarial and stochastic regret guarantees characteristic of FTRL. Hence, we achieve both statistical and computational efficiency.
@@ 1.1 Related Work
(i) Contextual Combinatorial Bandits. The contextual combinatorial bandit problem was introduced by Qin et al. (2014), who proposed the C2UCB algorithm for multi-item recommendations. This framework builds on earlier combinatorial semi-bandit models motivated by influence maximization in social networks Chen et al. (2013). Under the i.i.d. reward assumption, linear generalization across arms reduces the regret dependence on the action-set size $K$ from $\sqrt{K}$ to the context dimension $d$. Consequently, when $d\ll K$, the regret scales with $\sqrt{dT}$ instead of $\sqrt{KT}$, offering a substantial statistical advantage. Stochastic variants have explored Thompson sampling Wang and Chen (2018) and refined confidence sets Takemura et al. (2021); however, none of these algorithms provide optimal regret guarantees against an adaptive adversary.
@@
(ii) Semi-Bandit Feedback and Efficient Optimization. Combinatorial bandits were first formalized by Cesa-Bianchi and Lugosi (2012), who generalized Exp3 (Auer et al., 2002) from single arms to binary action vectors. Their algorithm, ComBand, operates under full-bandit feedback—the learner only observes an aggregated loss for the binary action vector they have chosen—and achieves $\widetilde{\mathcal{O}}\bigl(\sqrt{mKT\ln(K/m)}\bigr)$ regret. Follow-up work connected this update to mirror descent on combinatorial polytopes: ComponentHedge Koolen et al. (2010), CombEXP Combes et al. (2015), and the OSMD analysis of Audibert et al. (2014). Audibert et al. (2014) provide a concise taxonomy and prove matching lower bounds, showing that the $\sqrt{mKT}$ rate is information-theoretically optimal whenever semi-bandit feedback is available. These results are also summarized in (Bubeck and Cesa-Bianchi, 2012, Section 5.6.1). To reduce per-round runtime,  Neu and Bartók (2016) introduced a stochastic mirror-descent update that samples a single action and updates only the observed coordinates, thereby avoiding the full Bregman projection required by standard OSMD or FTRL. The trade-off is an extra logarithmic factor in the regret bound. Variance-reduced estimators and high-probability analyses later refined these guarantees (Zimmert et al., 2019), yet computing the exact Bregman projection still requires time linear in the number of arms $K$.
@@
(iii) Adversarial Linear Bandits and Best-of-Both-Worlds. When losses can adapt to the learner’s past, the i.i.d. assumption no longer holds. For single-arm actions ($m=1$) the classical EXP4 algorithm attains $\widetilde{\mathcal{O}}(\sqrt{T})$ regret, but its running time and memory scale with the number of experts—that is, the total number of deterministic policies mapping contexts to arms—which grows exponentially in the arm size $K$ (Auer et al., 2002).  Neu and Olkhovskaya (2020) mitigated this explosion with RealLinExp3, achieving $\widetilde{\mathcal{O}}(\sqrt{dKT})$ regret under the strong assumption that the context distribution is known. Liu et al. (2023) proposed an algorithm that achieves $\mathcal{O}((dm)^{3}\sqrt{T})$ regret bound when applied to combinatorial semi-bandit case with only access to $1$ context sample per time period $t$, though the resulting log-determinant FTRL remains intractable for the combinatorial bandits case because of exponentially many constraints Foster et al. (2020); Zimmert and Lattimore (2022) and the nonlinearity introduced by their lifting covariance technique. The first near-optimal treatment is the Matrix-Geometric-Resampling (MGR) algorithm of Zierahn et al. Zierahn et al. (2023), which achieves $\widetilde{\mathcal{O}}\bigl(\sqrt{mKT\max\{d,m/\lambda_{\text{min}}(\Sigma)\}}\bigr)$ regret but fails to provide best-of-both-worlds regret guarantees and relies on a sampling subroutine which requires $\mathcal{O}(\ln T)$ samples in each round. Similar to the issue of generalizing (Liu et al., 2023), a straightforward extension of the best-of-both-worlds (BOBW) algorithm from the linear contextual bandit setting, as studied by Kuroki et al. (2024), to the contextual combinatorial bandit framework yields an action space whose cardinality grows exponentially with the size of the combinatorial decision variables. This exponential growth renders the resulting optimization problem NP-hard and computationally infeasible for large-scale applications. More broadly, the BOBW question—achieving $\widetilde{\mathcal{O}}(\ln T)$ under stochastic contexts and $\widetilde{\mathcal{O}}(\sqrt{T})$ under adversarial ones without prior knowledge—originated in the $K$-armed bandit setting (Bubeck and Slivkins, 2012; Seldin and Slivkins, 2014) and has since been settled for linear bandits via data-dependent stability (Lee et al., 2021), and a streamlined FTRL scheme (Kong et al., 2023) and for contextual bandits via Exp4-style expert reductions (Pacchiano et al., 2022; Dann et al., 2023), at the cost of exponential policy-enumeration. In the purely combinatorial semi-bandit setting, hybrid-regularizer methods (Zimmert et al., 2019; Ito, 2021) achieve the optimal BOBW rates but do not consider contextual information and the stemming difficulty of inverse-covariance matrix estimation. Consequently, no prior method simultaneously handles unknown covariance, large/infinite policy classes, and combinatorial action structure while maintaining BOBW regret bounds. We close this gap with a Shannon-entropy FTRL algorithm that requires neither policy enumeration nor known $\Sigma$, yet still achieves optimal BOBW rates on arbitrary combinatorial action sets.
Table 1: Comparison of regret bounds across different settings.
Our Paper Qin et al. (2014) Zierahn et al. (2023) Ito et al. (2022) Kong et al. (2023)
Feedback Semi-bandit Full-bandit Semi/Full-bandit Graph bandit Linear bandit
Adv. Regret $\widetilde{\mathcal{O}}\!\left(\mathrm{poly}(d,m,K)\sqrt{T}\right)$ N/A $\widetilde{\mathcal{O}}\!\left(\mathrm{poly}(d,m,K)\sqrt{T}\right)$ $\widetilde{\mathcal{O}}\!\left(\sqrt{\alpha T}\right)$ $\widetilde{\mathcal{O}}\!\left(\sqrt{T}\right)$
Stoch. Regret ${\mathcal{O}}\!\left(\mathrm{poly}(d,m,K)(\ln T)^{3}\right)$ $\widetilde{\mathcal{O}}\!\left(d\sqrt{mT}\right)$ N/A ${\mathcal{O}}\!\left(\frac{\alpha(\ln T)^{3}}{\Delta_{\min}}\right)$ ${\mathcal{O}}\!\left(\frac{(\ln T)^{2}}{\Delta_{\min}}\right)$
@@ 2 Best-of-Both-Worlds Algorithm for Contextual Combinatorial Semi-Bandits
We begin our analysis of contextual combinatorial bandits by describing the interaction protocol that the learner follows and the problem settings that the best-of-both-worlds framework unifies. Given $K$ base arms, $m$ maximum number of base arms allowed to pull per round, an action set $\mathcal{A}{\subseteq}\{A\in\{0,1\}^{K}:\sum_{k=1}^{K}(A)_{k}\leq m\}$, and a context space $\mathcal{X}\subset\mathbb{R}^{d}$, the interaction protocol for the contextual combinatorial bandit problem proceeds as follows.
@@
In each round $t=1,\dots,T$ the interaction protocol proceeds as follows. The environment first chooses the loss coefficients $\theta_{t,1},\ldots,\theta_{t,K}\in\mathbb{R}^{d}$, and draws an i.i.d. context $X_{t}\sim\mathcal{D}$. The learner observes $X_{t}$ and plays the vector $A_{t}\in\mathcal{A}$. The learner then observes the arm-wise losses $\ell_{t}(X_{t},k)=\langle X_{t},\theta_{t,k}\rangle$ for every $k$ such that $(A_{t})_{k}=1$ and suffers the total loss $\sum_{k=1}^{K}\ell_{t}(X_{t},k)(A_{t})_{k}$. Consistent with the prior work on linear contextual bandits such as (Kuroki et al., 2024), we adopt the following assumption throughout this paper.
@@
Under this assumption, we distinguish between the adversarial and stochastic regimes as follows. In the adversarial regime, the unknown loss coefficients $\{\theta_{t,k}\}_{k=1}^{K}$ may vary arbitrarily with respect to time but are independent of the learner’s action sequence. In the stochastic regime, the loss function takes the form $\ell_{t}(X_{t},k)=\langle X_{t},\theta_{k}\rangle+\varepsilon(X_{t},k)$ for all $k\in[K]$, where $\theta_{k}$ is also unknown but fixed throughout the $T$ interaction periods, and $\varepsilon(X_{t},k)$ is an independent zero-mean noise. There is an additional intermediate regime that interpolates between the adversarial and stochastic regime, referred to as corrupted stochastic regime (Zimmert and Seldin, 2021; Kuroki et al., 2024). In this case, the loss function associated with each base arm $k$ is defined by $\ell_{t}(X_{t},k)=\langle X_{t},\theta_{t,k}\rangle+\varepsilon(X_{t},k)$, where $\varepsilon(X_{t},k)$ is independent zero-mean noise. On the other hand, the coefficients $\theta_{t,k}$ are such that there exist fixed and unknown vectors $\theta_{1},\ldots,\theta_{K}$ and that satisfy $\sum_{t=1}^{T}\max_{A\in\mathcal{A}}\sum_{k=1}^{K}\|\theta_{t,k}-\theta_{k}\|_{2}(A)_{k}\leq C$ for a fixed corruption level constant $C>0$. Note that $C=0$ corresponds to the stochastic regime and $C=\Theta(T)$ corresponds to the adversarial regime with additional zero-mean noise. For any fixed context vector $x\in\mathcal{X}$, we define $u^{\star}:\mathcal{X}\to\mathcal{A}$ as the optimal context-dependent action map that achieves minimum loss in hindsight, that is,
@@
where the expectation is taken with respect to the action-selection distribution chosen by the learner and the sequence of random contexts and loss coefficients chosen by the environment. As noted in the introduction, it is well-established in the literature that the optimal regret bounds are $\widetilde{\mathcal{O}}(\ln T)$ in the stochastic regime and $\widetilde{\mathcal{O}}(\sqrt{T})$ in the adversarial regime. Algorithms that achieve both rates simultaneously and without prior knowledge of the environment’s nature are referred to as “best-of-both-worlds” (Bubeck and Slivkins, 2012; Seldin and Slivkins, 2014).
@@
4:  Compute $\bar{A}_{t}(X_{t})\in\operatorname*{argmin}_{a\in\operatorname*{conv}(\mathcal{A})}\sum_{s=1}^{t-1}\sum_{k=1}^{K}\langle X_{t},\tilde{\theta}_{s,k}\rangle a_{k}+\psi_{t}(a)$
@@
12:   return $\widehat{\Sigma}_{t,k}^{+}=(I+\sum_{n=1}^{M_{t}}C_{n,k})/2$ for all $k\in[K]$
13:  End Subroutine
14:  Compute $\tilde{\theta}_{t,k}=\widehat{\Sigma}_{t,k}^{+}X_{t}\ell_{t}(X_{t},k)(A_{t})_{k}$ for all $k\in[K]$